Find the indicated probability.
P(A):50%
P(B):50%
A(A orB)= ______
P(A and B)= 0%

Answer by ikleyn(52778)

.

If P(A) = 0.5,  P(B) = 0.5  and  P(A and B) = 0, then


    P(A or B) = P(A) + P(B) - P(A and B) = 0.5 + 0.5 - 0 = 1 = 100%.    ANSWER

Solved.